# Vector-Quantized Vision Foundation Models for Object-Centric Learning

### v3

@@ 4.1. Set up the Benchmark
contains synthetic videos, each with up to 20 daily objects dropping and bumping. COCO [^4] https://cocodataset.org is a recognized real-world image dataset, and we use its challenging panoptic segmentation. VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC is a real-world image dataset, and we use its instance segmentation. We also report results on real-world video dataset YTVIS version HQ [^6] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube. We choose Physion [^7] https://physion-benchmark.github.io for visual prediction and reasoning as it contains common object interactions, requiring algorithms to learn dynamics like support, roll and link, then to predict and reason about future scene states.
@@
Using higher resolution: 384$\times$384 (336)
resolution=384$\times$384 (336) COCO #slot=7
@@
4.2. Evaluate on Object Discovery
Object discovery task intuitively shows how well those slots separate different objects. We evaluate the byproduct object segmentation accuracy with Adjusted Rand Index [^10] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_sco
re.html
(ARI), ARIfg (foreground), mean Intersection-over-Union [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO).
With unsupervised pretrained VFMs for OCL encoding, i.e., DINO2 ViT (version s/14), our VVO is instantiated as VQDINO. As shown in Table 1, our method consistently improves object discovery performance across all types of OCL decoding. With a Transformer decoder Tfd for auto-regressive decoding, VVO significantly outperforms SLATE and STEVE across all datasets. With MLP Mlp for mixture decoding, VVO shows a smaller advantage over DINOSAUR but is still effective. With a conditional Diffusion model Dfz for diffusion decoding, VVO surpasses SlotDiffusion on most datasets.
With supervised pretrained VFMs for OCL encoding, i.e., SAM2 Hiera+FPN (version t/16), our VVO is instantiated as VQSAM. As shown in Table 2, VVO consistently boosts those baseline methods’ object discovery performance across all decoding types on most datasets.
As shown in Table 3, whether using a higher input resolution or different aggregators, VVO maintains its superiority over baselines. As shown in Table 4, VVO outperforms recent general improvers for OCL training, i.e., GDR and MSF, which do not support mixture decoding like us though. VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
4.3. Evaluate on Set Prediction
@@
Table 5. Set prediction performance on COCO (panoptic, #slot=7).
Set prediction task directly shows how much object information those slots grasp. We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023). We measure top1 accuracy of the classified class labels while we measure the R2 score [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.r2_score.html of the regressed bounding box coordinates.
As shown in Table 5, compared with baseline DINOSAUR, our VVO, i.e., VQDINO Mlp , obtains better object classification and bounding box regression. Thus, our models’ object representations are of higher quality.
@@ 4.4. Deploy to the Downstream
On visual prediction, we evaluate the prediction errors per time step measured in normalized MSE between regressed and extracted slots. As shown in Figure 5 upper, VVO accumulates prediction errors slower than the baseline.
On visual reasoning, we evaluate the reasoning accuracies per time step between classified and ground-truth labels. As shown in Figure 5 lower, VVO accuracies are slightly lower at the beginning but much higher later than the baseline.
@@ 4.5. Ablate the Architecture
As shown in Table 6, VVO’s design of shared VAE and OCL encoder consistently outperforms separate VAE and OCL encoders, even when the latter employs another VFM for VAE encoding. Thus, VVO’s design of shared VFM representation quantization is crucial. However, the design of separate encoders remains prevalent among those baselines.
As shown in Table 7, VVO’s improved quantizer variant, based on tricks of Gumbel noises defined in Equation 11, annealing residual connection defined in Equation 12 and normalizing regularization defined in Equation 13, is superior to the naive VQ. The detailed effects of those tricks are shown in Figure 6. Adding Gumbel noises increases codebook utilization, contributing to more effective codes; Annealing residual connection improves VAE pretraining, contributing to smaller VAE reconstruction error.
@@
Table 6. VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders. Results are on COCO panoptic.
@@
Table 7. VVO’s VQ variant: All our three tricks are beneficial to the overall performance boosts. In comparison to VVO’s key designs, these tricks are more like the cherry on top. Results are on COCO panoptic with settings consistent with the above.
@@ 5. Analysis
We provide both mathematical modeling and statistical verification on why and how our VVO is better.
5.1. Aggregation as Clustering
Mathematic Modeling
@@ 5.1. Aggregation as Clustering
[…]
where $\bm{o}_{2}$ always contains $\bm{s}_{*}$, and is bounded by the separation plane between $\bm{o}_{1}$ and $\bm{o}_{2}$. The closer the boundary is to $\bm{s}_{*}$ the smaller $p_{2}$ is.
According to Figure 1 the first observation, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space. This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space. Therefore, $p_{2}$ in VFM space is bigger than in non-VFM space. $\square$
Statistical Verification
Equation 15 is very difficult to solve, but we can use Monte Carlo estimation:
(16) $p_{2}=\frac{1}{N}\Sigma_{i=1}^{N}\mathbbm{1}[\bm{v}_{i}\in\bm{o}_{2}]$
Specifically, we respectively use ResNet18, DINO2 ViT and SAM2 encoder to extract images into super-pixels, then apply KMeans clustering to find each clusters mean and std. From the Gaussian distributions defined by these mean and std, we repeatedly draw samples and calculate the ratio that such samples fall within their original clusters. This ratio is the correct aggregation probability.
Figure 8. Statistical correlation between correct aggregation probability and object discovery accuracy on VOC.
A step further, we can calculate the statistical correlation between the correct aggregation probability and the corresponding OCL performance. As shown in Figure 8, the correlation is roughly significant. Considering that correct aggregation probability can be calculated before the expensive OCL training, our Equation 15 is instructive to training better OCL models by selecting OCL encoders.
5.2. Shared Quantization and Optimization Noise
Mathematic Modeling
@@ 5.2. Shared Quantization and Optimization Noise
where $\frac{\partial f(\bm{Z})}{\partial\bm{Z}}$ is the Jacobian matrix of $f(\bm{Z})$ with respect to $\bm{Z}$. Anyway, this has similar structure to Equation 17 and thus does not alter our judgment above. $\square$
Statistical Verification
Figure 9. Gradient distributions tend to disperse or have outliers if not using our shared VFM feature quantization. Results here are collected at time step 10k, on dataset VOC.
The statistical bias in Equation 20 manifests as the gradient distribution of different layers, especially that of slots. No statistical bias is usually associated with a smooth zero-mean small-std Gaussian gradient distribution.
Our technique, shared VFM feature quantization as reconstruction targets, includes two key points, sharing and quantization. Thus we compare three cases: (i) Shared and quantized, where DINO features are both used for aggregation and quantized for reconstruction; (ii) Shared but not quantized, where DINO features are used for aggregation and direct for reconstruction without quantization; (iii) Not shared but quantized, where DINO features are used for aggregation while quantized SAM features are used for reconstruction.
As shown in Figure 9, using CE or MSE as the loss, our shared quantization achieves smoother Gaussian gradient distributions, which benefits OCL training. If not shared but quantized, then gradients tend to disperse; while if shared but not quantized, then gradients have some outliers.

### v5

@@ 4.1. Set up the Benchmark
contains synthetic videos, each with up to 20 daily objects dropping and bumping. COCO [^4] https://cocodataset.org is a recognized real-world image dataset, and we use its instance segmentation. VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC is a real-world image dataset, and we use its instance segmentation. We also report results on real-world video dataset YTVIS [^6] https://youtube-vos.org/dataset/vis version HQ [^7] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube. We choose Physion [^8] https://physion-benchmark.github.io for visual prediction and reasoning as it contains common object interactions, requiring algorithms to learn dynamics like support, roll and link, then to predict and reason about future scene states.
@@
Using higher resolution: $\times384$ (336)
[…]
@@
Table 5. Set prediction performance on COCO (#slot=7).
4.2. Evaluate on Object Discovery
Object discovery task intuitively shows how well those slots separate different objects. We evaluate all methods’ byproduct object segmentation accuracy with Adjusted Rand Index [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_sco
re.html
(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO) as metrics.
With unsupervised pretrained VFMs for OCL encoding, i.e., DINO2 ViT (version s/14), our VVO is instantiated as VQDINO. As shown in Table 1, our method consistently improves object discovery performance across all types of OCL decoding. With a Transformer decoder Tfd for auto-regressive decoding, VVO significantly outperforms SLATE and STEVE across all datasets. With a spatial broadcast MLP decoder Mlp for mixture-based decoding, VVO shows a smaller advantage over DINOSAUR but is still effective. With a conditional Diffusion model Dfz for diffusion-based decoding, VVO surpasses SlotDiffusion on most datasets.
With supervised pretrained VFMs for OCL encoding, i.e., SAM2 Hiera+FPN (version t/16), our VVO is instantiated as VQSAM. As shown in Table 2, VVO boosts all baselines’ object discovery performance across all decoding types on all datasets.
As shown in Table 3, whether using higher input resolution or different aggregators, VVO maintains its superiority over baselines. As shown in Table 4, VVO outperforms recent OCL general improvers, i.e., GDR and MSF. VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
4.3. Evaluate on Set Prediction
Set prediction task directly shows how much object information those slots grasp. We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023). We measure top1 accuracy [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.accuracy_score.ht
ml
accuracy of the classified class labels while measure R2 score [^14] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.r2_score.html of the regressed bounding box coordinates.
As shown in Table 5, compared with DINOSAUR, our VVO, i.e., VQDINO Mlp , obtains both better object classification and better object bounding box regression. Thus, our method extracts better slot representations for objects than the baseline.
@@ 4.4. Deploy to the Downstream
On visual prediction, we evaluate the per time step prediction errors measured in normalized Mean Squared Error (MSE) between regressed and extracted slots. As shown in Figure 5 upper, our VVO accumulates prediction errors slower than the baseline.
On visual reasoning, we evaluate the per time step reasoning accuracy between the classification outputs and ground-truth labels. As shown in Figure 5 lower, VVO’s accuracies are slightly lower at the beginning but much higher later than the baseline.
@@ 4.5. Ablate the Architecture
As shown in Table 6, VVO’s design of shared VAE and OCL encoder consistently outperforms separate VAE and OCL encoders, even when the latter employs another VFM for VAE encoding. Thus, VVO’s design of shared VFM representation quantization is superior to the prevalent design of separate VAE and OCL encoders.
As shown in Table 7, our improved quantizer variant for VVO, built upon tricks of Gumbel noises defined in Equation 11, annealing residual connection defined in Equation 12 and normalizing regularization defined in Equation 13, is superior to the naive VQ. The detailed effects of those tricks are shown in Figure 6. Adding Gumbel noises increases codebook utilization, contributing to more effective codes; Annealing residual connection improves VAE pretraining, contributing to smaller VAE reconstruction error.
@@
Table 6. VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders. Results are on COCO.
@@
Table 7. VVO’s VQ variant: All our three tricks are beneficial to the overall performance boosts. In comparison to VVO’s key designs, these tricks are more like the cherry on top. Results are on COCO with settings consistent with the above.
@@ 5. Analysis
We mathematically analyze our two key designs as below.
Aggregation as Clustering
@@
(15) $p_{2}=p(d_{s1}>d_{s2})=\int_{\bm{v}\in\bm{o}_{2}}\frac{1}{\sqrt{2\pi}\bm{\sigma}}e^{-\frac{1}{2}(\frac{\bm{v}-\bm{s}_{*}}{\bm{\sigma}})^{2}}d\bm{v}$
where $\bm{o}_{2}$ always contains $\bm{s}_{*}$, and is bounded by the separation hyper-plane between $\bm{o}_{1}$ and $\bm{o}_{2}$. The closer the boundary is to $\bm{s}_{*}$ the smaller the value of $p_{2}$ would be.
According to Figure 1 the first observation, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space. This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space. Thus, $p_{2}$ in VFM space is bigger.
Shared Quantization and Optimization Noise
@@
where $\frac{\partial f(\bm{Z})}{\partial\bm{Z}}$ is the Jacobian matrix of $f(\bm{Z})$ with respect to $\bm{Z}$. Anyway, this has similar structure to Equation 16 and thus does not alter our judgment above.
@@ 6. Conclusion
Acknowledgements.
We acknowledge the support of Finnish Center for Artificial Intelligence (FCAI), Research Council of Finland flagship program. We thank the Research Council of Finland for funding the projects ADEREHA (grant no. 353198), BERMUDA (362407) and PROFI7 (352788). We also appreciate CSC - IT Center for Science, Finland, for granting access to supercomputers Mahti and Puhti, as well as LUMI, owned by the European High Performance Computing Joint Undertaking (EuroHPC JU) and hosted by CSC Finland in collaboration with the LUMI consortium. Furthermore, we acknowledge the computational resources provided by the Aalto Science-IT project through the Triton cluster.